carmelle:

If x = 10, y = 5, and z = 11, what is the tangent ratio for g°? A. 1/2 B. 2/1 C. 10/11 D. 5/11

carmelle:

1 attachment
Bear20:

THE ANSWER IS SPOON

carmelle:

Are you going to help with the question?

Bear20:

yes

Bear20:

THFK{O_BIHYCGbhvbkn]

Bear20:

thats the answer

Macho14:

dw, I got you

Macho14:

The triangle ACB is a right triangle with these features: x represents the opposite side to the angle g y represents the adjacent side to the angle g The tangent ratio on an angle in a right triangle is equal to the length of the leg opposite to the angle divided by the length of the leg adjacent to the angle.

Macho14:

\[tangent(g)=x/y=10/5=?\] then you have your answer

carmelle:

Thank you!! The final answer is 2/1.

Macho14:

yes, otherwise known as 2

Macho14:

np